Use an inequality to solve the problem. Be sure to show the inequality and all of your work.
One side of a triangle is 2 cm shorter than the base, x. The other side is 8 cm longer than the base. What lengths of the base will allow the perimeter of the triangle to be at least 30 cm?

Answer by mananth(16946)

base =x
one side = x-2
other side = x+8
..
the perimeter should be at least 30 if not more
x+x-2+x+8>=30
3x+6>=30
-6
3x-6+6>=36
3x>=24
/3
x>=8
..
m.ananth@hotmail.ca